Calculate the buoyant force due to the surrounding air on a man weighing 700 N . Assume his average density is the same as that of water. Suppose that the density of air is 1.20 kg/m3.
Answer #1

mass of man=700/10=70 kilograms

volume of man=mass/density=70/1000=0.07 kg/m3

buoyant force is given by

airg 0.07 1.20 100.84newtons a2r